Derive the wave equation for the z-component of the electric field of an electromagnetic
wave and show that the wave propagates in free space with the speed of light .
1
Expert's answer
2020-02-24T10:58:30-0500

Let the electric field vector is in the z direction, magnetic field vector is in the direction of the y axis. so the propagation of the wave is in the direction of the x direction, and electric and magnetic field vectors are in the function of x and t.

So, "\overrightarrow{E}(z,t)=E(z,t) \hat{k}"

Magnetic field "\overrightarrow{B}(y,t)=B(y,t)\hat{j}"

Now, as per the maxwell's equation for the free space

"\nabla . E=0" and "\nabla.B=0"

"\nabla\times E=-\dfrac{\partial B}{\partial t}"

and "\nabla\times B=\mu_o\epsilon_o\dfrac{\partial E}{\partial t}"

Now, taking the curl of the electric field vector



So, "-\dfrac{\partial B}{\partial t}=\dfrac{\partial E}{\partial z}" ------(i)

"-\dfrac{\partial B}{\partial z}=-\mu_o\epsilon_o\dfrac{\partial E}{\partial z}"

Similarly for the magnetic field


now taking the partial derivative of the equation (i)

"\dfrac{\partial^2 E}{\partial z^2}=-\dfrac{\partial (\partial B)}{\partial t\partial z}=\mu_o\epsilon_o\dfrac{\partial^2 E}{\partial z^2}" -------(ii)

We know that the general wave equation for the traveling wave.

"\dfrac{\partial^2\psi}{\partial x^2}=\dfrac{1}{\nu}\dfrac{\partial^2\psi}{\partial t^2}" ------(iii)

From the equation (ii) and (iii)

"c=\dfrac{1}{\sqrt{\mu_o\epsilon_o}}=2.997\times 10^8m/sec"